You(mass 60kg) and your friend (mass 80kg) are on the enterprise at valley fair. You each experience 1,200N of force during the ride. How many g’s is that for you and for your friend
1
Expert's answer
2020-11-25T07:12:03-0500

My accleration ( due to second Newton's law) is:


"a_1 = \dfrac{F}{m_1} = \dfrac{1200N}{60kg} = 20m/s^2"

Since "g \approx 10m/s^2", the acceleration is 2g's.

My friend's acceleration is:


"a_2 = \dfrac{F}{m_2} = \dfrac{1200N}{80kg} = 15m/s^2"

This is 1.5g's.

Answer. 2 and 1.5 g's.